# Cross-Layer Vision Smoothing: Enhancing Visual Understanding via Sustained Focus on Key Objects in Large Vision-Language Models

### v1

@@ 4.1 Setup
  • AMBER (Wang et al. 2023): a human-annotated benchmark including three types of hallucination evaluation questions: existence, attribute, and relation.
  • R-Bench (Wu et al. 2024): a benchmark for evaluating relation hallucination, which pose a greater challenge to the model’s visual understanding ability. This benchmark provides multiple randomly constructed subsets with a balanced number of Yes and No questions, and we use the subset with ID 1. We conduct experiments using image-level questions.
We also conduct experiments on the following benchmarks to further validate our proposed CLVS method:
  • POPE (Li et al. 2023): a benchmark for evaluating existence hallucination. It consists of three subsets: Random, Popular, and Adversarial, each constructing negative examples in a distinct manner.
  • MME (Fu et al. 2024): a comprehensive evaluation benchmark. We use perception tasks to test the model’s visual understanding ability following previous works, which includes four types of questions: Existence, Color, Count, and Position.
@@
  • VCD (Leng et al. 2024): a method that stimulates visual uncertainty by adding noise to images and purifies visual information using contrastive decoding.
  • PAI (Liu, Zheng, and Chen 2024): a method that directly amplifies visual attention and performs contrastive decoding with pure textual context to purify visual information.
  • VAR (Kang et al. 2025): a method that identifies visual sink tokens and redistributes the model’s attention from sink tokens to other visual tokens.
  • AD-HH (Yang et al. 2025): a method that identifies attention heads that cause hallucinations. When the attention of these heads to text exceeds a threshold, their attention to text is set to zero, thereby enhancing visual attention.
  • VAF (Yin, Si, and Wang 2025a): a method that redistributes the model’s attention from system tokens to visual tokens in the middle layers where the model performs visual understanding.
  • TARAC (Xie et al. 2025): a method that also optimizes visual attention in the middle layers. It accumulates visual attention from each generation step and injects it into the visual attention of the current step.
The implementation details of baselines are provided in the Appendix A.
@@
In all experiments, LVLMs use greedy search for next-token prediction. We set $\beta=0.8$ in Eq. (5) for all LVLMs. The value of $\gamma$ in Eq. (7) is set to 0.8 for LLaVA-1.5, 0.7 for LLaVA-Next and Qwen2.5-VL. The setting of uncertainty threshold $\delta$ is based on the statistical analysis (see Figure 4), with a value of 0.5. All models are tested at the 7B or 8B parameter scale unless explicitly stated otherwise. We conduct our experiments on a single NVIDIA A800 40G GPU.
Method Acc. Prec. Rec. F1 Acc. Prec. Rec. F1
LLaVA-1.5-7B LLaVA-1.5-13B
Vanilla 67.8 61.6 95.3 74.8 68.2 61.7 96.9 75.4
VCD (Leng et al. 2024) 68.2 62.1 94.6 75.0 68.9 62.5 95.2 75.5
PAI  (Liu, Zheng, and Chen 2024) 68.7 62.4 95.3 75.4 71.0 64.2 95.9 76.9
VAR (Kang et al. 2025) 68.0 61.8 95.2 75.0 68.0 61.5 97.0 75.3
AD-HH (Yang et al. 2025) 67.6 61.5 95.4 74.8 68.6 62.1 96.7 75.6
VAF (Yin, Si, and Wang 2025a) 66.8 60.7 96.2 74.5 67.7 61.3 96.9 75.1
TARAC (Xie et al. 2025) 67.7 61.5 95.5 74.8 68.8 62.2 96.7 75.7
CLVS 69.0 62.5 95.7 75.6 71.3 64.6 95.2 77.0
LLaVA-Next Qwen2.5-VL
Vanilla 81.9 80.1 85.3 82.6 81.6 80.2 84.2 82.1
VCD (Leng et al. 2024) 80.2 78.3 83.8 81.0 79.5 78.3 82.1 80.1
VAF (Yin, Si, and Wang 2025a) 81.3 78.2 87.0 82.4 81.5 80.7 83.1 81.9
TARAC (Xie et al. 2025) 81.8 80.3 84.5 82.3 81.6 80.9 83.1 82.0
CLVS 81.9 79.3 86.6 82.8 81.8 80.4 84.2 82.3
Table 2: Results on R-Bench. “Prec.” and “Rec.” denote Precision and Recall, respectively.
Method Accuracy Recall F1
LLaVA-Next
Vanilla 87.39 77.09 85.94
CLVS 88.49 (+1.10) 79.78 (+2.69) 87.39 (+1.45)
Qwen2.5-VL
Vanilla 89.91 82.87 89.15
CLVS 90.16 (+0.25) 83.07 (+0.20) 89.40 (+0.25)
Table 3: Results on the Random subset in POPE.
4.2 Main Results
@@
Table 1 shows the results on AMBER, a comprehensive hallucination evaluation benchmark that includes tests for three major types of hallucinations. Our method achieves notable improvements over baselines on multiple LVLMs, particularly in understanding object relationships. Compared with other tasks, the improvement in determining object existence is relatively small, especially when using models such as Qwen2.5-VL and LLaVA-Next, which possess stronger visual comprehension capabilities. This is because our approach leverages the inherent attention advantages of LVLMs to maintain focus on key objects, thereby substantially enhancing performance in complex relational reasoning. However, in object-existence determination tasks where relevant objects may be absent from the image, our method cannot establish focused attention on key objects and thus provides limited improvements.
@@
Table 2 shows the results on the R-Bench benchmark. Our method achieves significant improvements in evaluating relational hallucinations, attaining state-of-the-art F1 scores across all LVLMs, including strong models such as Qwen2.5-VL and LLaVA-Next. These results are consistent with those on AMBER, further demonstrating the versatility of our method.
4.3 Results on POPE
As discussed earlier, our method provides limited improvements when key objects are absent from the image. For example, when the answer to the question “Is there a/an {object} in the image?” is “False”, there are no relevant key objects to attend to, which limits the effectiveness of our approach. To further examine this issue, we conduct additional experiments on the POPE dataset, a benchmark designed for existence hallucination. The results are presented in Table 3. We find that while our method does not yield substantial gains in accuracy, it achieves a notable increase in recall. This indicates that when key objects are present, our method effectively enhances the model’s visual understanding by sustaining attention on them.
4.4 Results on MME
Figure 5: Radar chart of results on the MME benchmark.
We further evaluate our method on a broader range of tasks using the MME benchmark, with results shown in Figure 5. Our method achieves the highest overall score among all baselines except on the position task. We conjecture that the position task relies heavily on spatial information in the image, while our approach primarily prolongs LVLMs’ focus on key objects and therefore does not enhance their understanding of spatial relationships. Nevertheless, our method yields significant improvements on the color and count tasks, where key objects are present.
Method Acc. Prec. Rec. F1
CLVS 75.0 93.3 67.1 78.1
w/o Early Termination 74.9 93.3 67.0 78.0
w/o Unified Position 74.8 93.1 67.0 77.9
@@ 4.4 Results on MME
Table 4: Ablation Study on AMBER with LLaVA-1.5-7B. “w/o Early Termination” represents performing vision smoothing until the final layer. “w/o Unified Position” represents employ original position indices in the first layer.
@@ 4.5 Ablation Study
We conduct an ablation study on CLVS, with results presented in Table 4. We observe positive contributions from both unified positional indices and early termination. For unified positional indices, the improvements are considerably smaller than those achieved by vision smoothing, since this operation is applied only in the first layer and its influence gradually diminishes during subsequent smoothing. Therefore, simply increasing the amount of information attended to by altering positional indices does not directly enhance visual understanding. Regarding early smoothing termination, we confirm that it has no negative impact on model performance, consistent with previous findings that visual understanding in LVLMs primarily occurs in the early and middle layers. Thus, early termination reduces computational cost without sacrificing performance. As a result, we confirm the effectiveness of all elements in our CLVS.
4.6 Analysis
Key Object Attention Analysis.
To validate the effectiveness of CLVS in strengthening advantageous visual attention, we conduct experiments on 500 positive samples from POPE and use manually annotated instance segmentations from the MSCOCO dataset to localize key objects. We measure the attention to a key object in layer $l$ using $a_{\textrm{obj}}^{(l)}=\sum_{i=1}^{N_{v}}\bm{a}^{(l)}[i]\cdot\frac{\mathrm{Area}(\textrm{patch}_{i}\cap\textrm{obj})}{\mathrm{Area}(\textrm{patch}_{i})}$, where $N_{v}$ is the number of visual tokens and $\bm{a}^{(l)}[i]$ is the average attention value for visual token $i$ across attention heads. We analyze the peak-layer distribution $l_{p}$ ($l_{p}=\mathrm{argmax}_{l}a_{\textrm{obj}}^{(l)}$) and the relative increase of $a_{\textrm{obj}}^{(l)}$ in CLVS over vanilla method (clipped at 100%). The statistical results are shown in Figure 6. In the vanilla method, attention to key objects typically emerges around layer 14 but quickly diminishes, consistent with prior findings (Zhang et al. 2025). In contrast, CLVS sustains focus on key objects, as evidenced by a large relative increase in attention weights in the lower layers, with an average gain of 31.65% from layers 15 to 17. Since early smoothing termination is applied starting in the middle layers, the model no longer preferentially attends to key objects above these layers. This enhanced attention facilitates better understanding of object-related information, enabling CLVS to reduce attribute and relation hallucinations.
@@ 4.6 Analysis
$\beta$ and $\gamma$ are two important hyperparameters that directly affect the performance of our method. We conduct experiments by varying their values, with results shown in Figure 7. We find a notable drop in performance when $\beta$ approaches 1 (with $\gamma$ fixed at 0.8), as the vision memory plays a smaller role in smoothing the visual attention distribution. We also observe a performance drop when $\gamma$ is small (with $\beta$ fixed at 0.8), since the vision memory may update too frequently. Very large or very small values of $\beta$ cause dramatic changes in the attention distributions across layers. These results are consistent with our hypothesis that sustained focus on key objects is essential.
@@ 5 Conclusion
Acknowledgments
This work is supported by Zhongguancun Academy Project No.20240103.
Appendix A Implementation Details for Baselines
  • VCD (Leng et al. 2024): We add 500 steps of Gaussian noise to the original image to construct the contrastive visual input, and the hyperparameters for contrastive decoding are set to $\alpha=1$, $\beta=0.1$.
  • PAI (Liu, Zheng, and Chen 2024): We activate the PAI method starting from layer 2 (counting from 0), with hyperparameters set to $\alpha=0.5$, $\gamma=1.1$.
  • VAR (Kang et al. 2025): The hyperparameters in this method are set as follows: $\tau=20$, $\rho=0.5$, and $p=0.6$.
  • AD-HH (Yang et al. 2025): We set the hallucination threshold $\tau$ for attention heads to 0.5. For the definition of hallucination heads, please refer to the original paper.
  • VAF (Yin, Si, and Wang 2025a): We set $\beta=0.1$, $\alpha=0.15$, and activate this method from layer 9 to layer 14 (counting from 0).
  • TARAC (Xie et al. 2025): We set $\beta=0.5$, $\alpha=0.5$, and activate this method from layer 9 to layer 15 (counting from 0).
@@ Appendix A Implementation Details for Baselines
CLVS 195.0 163.3 123.3 170.0 140.0 791.7 190.0 168.3 115.0 175.0 147.5 795.8
Table 5: The results on MME. Total denotes the overall score.
Model Method Random Popular Adversarial
Acc. Rec. F1 Acc. Rec. F1 Acc. Rec. F1
LLaVA-v1.5 Vanilla 88.3 81.6 87.5 85.0 81.6 84.5 81.1 81.6 81.2
CLVS 89.0(+0.7) 84.5(+2.9) 88.5(+1.0) 84.1(-1.0) 84.5(+2.8) 84.2(-0.3) 80.1(-1.0) 84.4(+2.8) 81.0(-0.2)
LLaVA-Next Vanilla 87.4 77.1 85.9 85.5 77.1 84.2 82.3 77.1 81.4
CLVS 88.5(+1.1) 79.8(+2.7) 87.4(+1.5) 86.1(+0.7) 79.8(+2.7) 85.2(+1.1) 82.7(+0.4) 79.8(+2.7) 82.2(+0.8)
Qwen2.5-VL Vanilla 89.9 82.9 89.2 87.9 82.9 87.3 84.8 82.9 84.5
CLVS 90.2(+0.3) 83.1(+0.2) 89.4(+0.3) 87.6(-0.3) 83.1(+0.2) 87.1(-0.2) 84.7(-0.1) 83.1(+0.2) 84.5(-0.1)
Table 6: The results on POPE with average scores among MSCOCO, A-OKVQA, and GQA.
Appendix B More Experimental Results
B.1 Results on MME

### v2

@@ 4.1 Setup
  • AMBER [AMBER]: a human-annotated benchmark that includes both visual question answering (VQA) and image captioning tasks. The VQA task encompasses three types of hallucination: existence, attribute, and relation. The image captioning task evaluates object hallucination.
  • R-Bench [R-Bench]: a benchmark for evaluating relation hallucination, which pose a greater challenge to the model’s visual understanding ability. This benchmark provides multiple randomly constructed subsets with a balanced number of Yes and No questions, and we use the subset with ID 1. We conduct experiments using image-level questions.
  • POPE [POPE]: a benchmark for evaluating existence hallucination. It consists of three subsets: Random, Popular, and Adversarial, each constructing negative examples in a distinct manner.
  • CHAIR [chair]: a benchmark for evaluating object hallucination in image captioning, assessing captions by annotations from the MSCOCO dataset.
@@
  • VCD [VCD]: a method that stimulates visual uncertainty by adding noise to images and purifies visual information using contrastive decoding.
  • PAI [PAI]: a method that directly amplifies visual attention and performs contrastive decoding with pure textual context to purify visual information. To evaluate the effect of enhanced visual attention, we denote the variant without contrastive decoding as PAI and the variant with contrastive decoding as PAICD.
  • VAR [VAR]: a method that identifies visual sink tokens and redistributes the model’s attention from sink tokens to other visual tokens.
  • AD-HH [AD-HH]: a method that identifies attention heads that cause hallucinations. When the attention of these heads to text exceeds a threshold, their attention to text is set to zero, thereby enhancing visual attention.
  • TARAC [TARAC]: a method that optimizes visual attention in the middle layers. It accumulates visual attention from each generation step and injects it into the visual attention of the current step.
@@
In all experiments, LVLMs use greedy search for next-token prediction. We set $\beta=0.8$ in Eq. 5 and $\gamma=0.8$ in Eq. 7 for all LVLMs. The uncertainty threshold $\delta$ is set to 0.5 for LLaVA-7B and LLaVA-13B, 0.9 for LLaVA-Next. We conduct our experiments on a single NVIDIA A800 40G GPU.
4.2 Results on Visual Understanding Task
@@
Table 1 shows the results on AMBER, a comprehensive hallucination evaluation benchmark that includes tests for three major types of hallucinations. A series of baselines based on visual attention optimization achieve significant improvements over the Vanilla results, highlighting the importance of visual attention in enabling LVLMs to understand visual content. PAICD achieves the second-best performance across all experiments; however, it requires two forward passes for contrastive decoding, doubling the inference cost. In contrast, our method achieves better performance with only a single forward pass, introducing no significant increase in computational overhead. Additionally, we observe that CLVS does not achieve the best performance on the existence hallucination task in the LLaVA-Next experiments. This is because the base model’s performance is already close to the upper bound of the dataset, and we notice some annotation noise in the data—making it difficult to determine ground truth accurately—which may lead to performance fluctuations. To further evaluate CLVS on existence hallucination, we additionally select the POPE benchmark for evaluation.
Results on POPE.
We further evaluate CLVS on existence hallucination, and the results are shown in the left part of Tab. 2. Our method achieves consistent and significant improvements across all three base models, and outperforms all other strategies under various negative sample selection schemes, demonstrating its effectiveness in enhancing the model’s perceptual awareness of objects. Additionally, we observe that CLVS significantly outperforms other visual attention optimization methods on existence hallucination, which strongly validates the importance of sustained visual focus for robust visual understanding.
@@
In the AMBER benchmark, CLVS shows strong performance in understanding object relationships. To further evaluate its effectiveness in enhancing model comprehension of object relations, we conduct experiments on the R-Bench benchmark, with results presented in the right part of Tab. 2. CLVS achieves the overall best performance in this Benchmark. Although it is slightly lower than PAICD on LLaVA-13B, it still surpasses PAI, demonstrating that our method retains a clear advantage in enhancing the model’s visual understanding capabilities.
4.3 Results on Image Captioning Task
CLVS enhances visual understanding by extending the model’s visual attention, and its effectiveness is thoroughly validated across multiple visual understanding benchmarks. Unlike visual understanding tasks, image captioning requires not only accurate visual perception but also coherent integration of textual features. To evaluate CLVS’s performance on image captioning, we conduct experiments on two dedicated captioning benchmarks. The results are presented in Tab. 3.
CLVS significantly reduces hallucinations in image captioning, achieving substantial improvements over the Vanilla baseline on both captioning benchmarks and matching the performance of state-of-the-art methods.
The reason CLVS achieves suboptimal performance in image captioning is that, during the process of generating descriptions, not every token requires visual information. When generating subwords, punctuation, or function words—tokens with strong language priors—the model does not rely on visual attention. In such cases, extending visual attention provides limited benefit. Despite this, sustained attention remains beneficial when the model generates content words, leading to the significant improvements achieved by CLVS.
@@ 4.3 Results on Image Captioning Task
Table 4: Ablation study on AMBER with LLaVA-7B. “w/o Early Termination” represents performing vision smoothing until the final layer. “w/o Unified Position” represents employing the original position indices in the first layer. “TTFT” represents “Time to First Token”. The AMBER metric is calculated as $(1-\mathrm{CHAIR}+\mathrm{F1})/2$.
Setting CHAIR Cover Hal Cog Acc. Prec. Rec. F1 AMBER TTFT
Vanilla 7.1 50.7 32.5 3.8 72.1 92.2 63.0 74.8 34.35 $\times 1.000$
CLVS 4.2 48.0 21.9 2.1 76.2 92.5 69.9 79.6 38.20 $\times 1.035$
w/o Early Termination 5.4 47.7 24.1 1.9 76.3 92.4 70.0 79.7 37.65 $\times 1.111$
w/o Unified Position 4.2 48.3 22.9 1.7 74.6 93.4 66.4 77.6 37.20 $\times 1.031$
4.4 Key Object Attention Analysis.
To validate the effectiveness of CLVS in strengthening advantageous visual attention, we conduct experiments on 500 positive samples from POPE and use manually annotated instance segmentations from the MSCOCO dataset to localize key objects. We measure the attention to a key object in layer $l$ as follows:
$a_{\textrm{obj}}^{(l)}=\sum_{i=1}^{N_{v}}\bm{a}^{(l)}[i]\cdot\frac{\mathrm{Area}(\textrm{patch}_{i}\cap\textrm{obj})}{\mathrm{Area}(\textrm{patch}_{i})}$ (9)
where $N_{v}$ is the number of visual tokens and $\bm{a}^{(l)}[i]$ is the average attention value for visual token $i$ across attention heads. We analyze the peak-layer distribution $l_{p}$ ($l_{p}=\mathrm{argmax}_{l}a_{\textrm{obj}}^{(l)}$) and the relative increase of $a_{\textrm{obj}}^{(l)}$ in CLVS over vanilla method (clipped at 100%). The statistical results are shown in Fig. 5. In the vanilla method, attention to key objects typically emerges around layer 14 but quickly diminishes, consistent with prior findings [ViCrop]. In contrast, CLVS sustains focus on key objects, as evidenced by a large relative increase in attention weights in the lower layers, with an average gain of 31.65% from layers 15 to 17. Since early smoothing termination is applied starting in the middle layers, the model no longer preferentially attends to key objects above these layers. This enhanced attention facilitates better understanding of object-related information, enabling CLVS to reduce attribute and relation hallucinations.
@@ 4.5 Ablation Study
We conduct an ablation study on AMBER, with results presented in Tab. 4.
Unified Position.
We uniform the position ids in visual tokens in the first layer to prevent position bias in visual attention. This method does not show significant effect in image captioning, as the task typically involves longer textual contexts, which dilute the impact of visual position bias compared to visual understanding tasks. Nevertheless, removing this module still leads to a slight increase in hallucination rate. In contrast, for visual understanding tasks, visual position bias can distort the attention distribution over the visual regions, thereby undermining the effectiveness of sustained visual attention.
Early Termination.
LVLMs primarily process visual information in the early and middle layers; therefore, we terminate CLVS when visual understanding completion to reduce unnecessary computation. The TTFT results confirm that this mechanism accelerates inference speed while maintaining visual understanding performance with negligible degradation. Moreover, we observe that Early Termination yields particularly notable benefits in image captioning. This is because, when generating tokens with strong language priors, the model relies less on visual input [CICD] and reaches convergence earlier [IBD]. In such cases, prolonging visual attention inappropriately can lead to adverse effects and increase hallucinations. Early Termination mitigates this issue by timely deactivating sustained focus, thereby improving both efficiency and generation accuracy.
@@
$\beta$ and $\gamma$ are two important hyperparameters that directly affect the performance of our method. We conduct experiments by varying their values, with results shown in Fig. 6. We find a notable drop in performance when $\beta$ approaches 1 (with $\gamma$ fixed at 0.8), as the vision memory plays a smaller role in smoothing the visual attention distribution. We also observe a performance drop when $\gamma$ is small (with $\beta$ fixed at 0.8), since the vision memory may update too frequently. Very large or very small values of $\beta$ cause dramatic changes in the attention distributions across layers. These results are consistent with our hypothesis that sustained focus on key objects is essential.
@@ 5 Conclusion
[…]
Supplementary Material
Table 5: The results on MME. Total denotes the overall score.
LLaVA-1.5-7B LLaVA-1.5-13B
@@
CLVS 195.0 158.3 123.3 175.0 132.5 784.1 195.0 163.3 115.0 180.0 140.0 793.3
Table 6: Results on AMBER. “All” denotes the aggregated results over all test samples.
Model Method Existence Attribute Relation All
Acc. $\uparrow$ F1 $\uparrow$ Acc. $\uparrow$ F1 $\uparrow$ Acc. $\uparrow$ F1 $\uparrow$ Acc. $\uparrow$ F1 $\uparrow$
Qwen-Chat Vanilla 95.7 97.8 81.8 83.0 55.5 64.5 83.6 88.3
CLVS 95.9 97.9 82.2 83.2 57.5 65.1 84.0 88.5
Qwen2.5-VL Vanilla 95.3 97.5 83.5 84.4 73.9 75.8 86.5 90.1
CLVS 95.0 97.4 83.9 84.6 74.5 76.3 86.6 90.2
A Detailed Experimental Setup
To assess the effectiveness of our method in alleviating multimodal hallucination, we conduct evaluations on four widely used benchmarks: one generative benchmark (CHAIR [chair]), two discriminative benchmark (POPE [POPE], R-Bench [R-Bench]), and one hybrid benchmark (AMBER [AMBER]).
CHAIR.
CHAIR measures the proportion of hallucinated objects—that is, objects generated by the model but absent from the ground-truth annotations. Following prior work, we randomly sample 500 images from the MSCOCO dataset. This benchmark reports two metrics, CHAIRs and CHAIRi, defined as:
CHAIRs $=\frac{|\text{Hallucinated Objects}|}{|\text{All Objects}|},$ (10)
CHAIRi $=\frac{|\text{Hallucinated Captions}|}{|\text{All Captions}|}.$
POPE.
POPE is a widely used benchmark for measuring object hallucination by prompting LVLMs to judge whether a specific object appears in a given image. It contains three datasets—MSCOCO, A-OKVQA, and GQA—each evaluated under three negative sampling strategies: Random, Popular, and Adversarial. Each subset consists of 3,000 questions paired with 500 images. Accuracy and F1 score are used as the primary metrics.
AMBER.
AMBER integrates both generative and discriminative assessments and is conducted on a curated set of 1,004 images. Besides image captioning, the benchmark includes 14,216 questions evaluating hallucinations in object, attribute, and relation understanding. AMBER provides multiple metrics—CHAIR, Cover, Hal, and Cog. Given the annotated object set $A_{obj}={obj_{1}^{A},obj_{2}^{A},\cdots,obj_{n}^{A}}$ and the generated object set $R^{\prime}{obj}$, these metrics are computed as:
CHAIR $=1-\frac{\mathrm{len}(R^{\prime}{obj}\cap A_{obj})}{\mathrm{len}(R^{\prime}{obj})},$ (11)
Cover $=\frac{\mathrm{len}(R^{\prime}{obj}\cap A_{obj})}{\mathrm{len}(A_{obj})},$
Hal $=\frac{{\text{CHAIR}>0}}{{\text{All Caps}}},$
Cog $=\frac{\mathrm{len}(R^{\prime}{obj}\cap H{obj})}{\mathrm{len}(R^{\prime}{obj})},$
where HobjH{obj} denotes the set of hallucinated target objects predicted by the LVLMs, and All Caps refers to all generated captions.
R-Bench.
R-Bench is a benchmark for evaluating relation hallucination, which pose a greater challenge to the model’s visual understanding ability. This benchmark provides multiple randomly constructed subsets with a balanced number of positive and negative samples. We use the subset with ID 1, which contains 5,530 samples. We conduct experiments using image-level questions.
A.1 Implementation Details for Baselines
  • VCD [VCD]: We add 500 steps of Gaussian noise to the original image to construct the contrastive visual input, and the hyperparameters for contrastive decoding are set to $\alpha=1$, $\beta=0.1$.
  • PAI [PAI]: We activate the PAI method starting from layer 2 (counting from 0), with hyperparameters set to $\alpha=0.5$, $\gamma=1.1$.
  • AD-HH [AD-HH]: We set the hallucination threshold $\tau$ for attention heads to 0.5. For the definition of hallucination heads, please refer to the original paper.
  • VAF [VAF]: We set $\beta=0.1$, $\alpha=0.15$, and activate this method from layer 9 to layer 14 (counting from 0).
  • TARAC [TARAC]: We set $\beta=0.5$, $\alpha=0.5$, and activate this method from layer 9 to layer 15 (counting from 0).
B Results on MME